OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y =square root of x (9,3)

OpenStudy (jonnyvonny):

First, you get the derivative of x^.5. Then, you plug in x=9 into the derivative to get you the instantaneous rate of change at x=9. Then, plug that slope, along with the point (9,3), into y=mx+b and solve for b. Then, still use the instantaneous slope that you found, and plug that in, along with the b you found, into y=mx+b Lastly, be happy you (hopefully) got the right answer. Any questions?

OpenStudy (anonymous):

why x^5? @JonnyVonny

OpenStudy (jonnyvonny):

x^.5

OpenStudy (jonnyvonny):

That is the square root of x.

OpenStudy (anonymous):

so the derivitive is 1/2x

OpenStudy (anonymous):

@JonnyVonny

OpenStudy (jonnyvonny):

No, sorry, 1 sec.

OpenStudy (jonnyvonny):

Okay, so, you have to do the chain rule to take the derivative of (x)^.5

OpenStudy (anonymous):

is it 1/2x @JonnyVonny

OpenStudy (jonnyvonny):

No, it is .5x^-.5\[\frac{ 1 }{ 2 \sqrt{x} }\]

OpenStudy (jonnyvonny):

Do you see how?

OpenStudy (anonymous):

i got y=1/6x+3/2